solve for x. (answer I got was either Xsquared-2x-6x+12 or zero?)

...____4____...= 1 -..._1_
.... X2+8X+12........... X+6

Answer by mukhopadhyay(490)

____4____...= 1 -..._1_
X2+8X+12........... X+6
4/(x+2)(x+6) = (x+5)/(x+6)
=> (x^2+7x+10) = 4
=> (x+6)(x+1) = 0
=> x=-6 or x=-1